Currently Nathan deposits $300 at the end of each month into an IRA and his company will match 40% of his deposit amount. He will retire in 45 years. Assuming his account will earn 8.5% interest rate (APR), how much he can withdraw monthly after his retirement after-tax basis? (Assume he will live for another 25 years after retirement, his average tax rate will be 20%, and his deposit amount will remain constant).
Answer #1

Nathan deposits =

300

Employer contribution = 40% of 300=

120

So, monthly deposit in IRA total at end of the month = 300+120=

420

Time 45 years or 45*12=

540 months

Interest rate 8.5% compounded monthly

So, Monthly interest rate 8.5%/12

0.708333%

Future value of annuity (F) due at end of the period = P * { (1+r)^n - 1 } / r

420 * ((1+0.708333%)^540 - 1)/0.708333%

2621921.437

So, after 45 years, value of investments shall be $2621921.44

Afer 45 years, Annuity investment value (P)=

2621921.437
interest rate monthly (r)= 0.708333%

withdrawl amount for 25 years = 25*12=

300 months

Withdrawl amount before tax basis shall be calculated by annuity formula =

Annuity payment = P*r (1+r)^n / ((1+r)^n -1)

2621921.437 * 0.708333%*(1+0.70833%)^300/((1+0.708333%)^300-1)

21112.22578

After tax withdrawl = before tax *(1-t)

21112.22578*(1-20%)

16889.78062

So, after tax withdrawl can be made $16,889.78